OpenStudy (anonymous):

1. 3t+5=2 2. -16=6a-4 3. y\3+2x=10 4. 1+k\4=-9 5. 8=n\-7-5 6. 3-c=7 7. 2a-8a=24 8. 2n+5=27 9. 4n-3=-7 10. n\2-10=5 11. n\6+6=-3 1. 3t+5=2 2. -16=6a-4 3. y\3+2x=10 4. 1+k\4=-9 5. 8=n\-7-5 6. 3-c=7 7. 2a-8a=24 8. 2n+5=27 9. 4n-3=-7 10. n\2-10=5 11. n\6+6=-3 @Mathematics

jhonyy9 (jhonyy9):

1. t=-1 2. a=-2 3. y+6x=30 so y=30-6x 4. 4+k=-36 so k= -40 5. -56=n+35 so n= - 91 6. c=-4 7. a=-4 8. 2n=22 so n=11 9. 4n=-4 so n= -1 10 n-20=10 so n=30 11. n+36=-18 so n= - 54